You are working on the OSU audit for the year ended 6/30/2018.

You have been assigned to audit

football ticket revenue for the home football games.

There were six homes games during this fiscal

year; one non-conference game and five conference games. One of the conference games was played

against OU.

The football stadium has 58,000 seats available; allocated as follows:

Box

2,000 seats

Club Level

6,000 seats

All other

50,000 seats

Ticket prices for the non-conference game

Box

$150 per ticket

Club Level

$90 per ticket

All other

$50 per ticket

Attendance for the non-conference game was allocated as follows:

Box

1,000 tickets sold

Club Level

4,000 tickets sold

All other

40,000 tickets sold

Ticket prices for conference games were in the following ranges.

For all games except OU, the

amounts charged were the lowest prices. For OU, the maximum ticket price was charged for all tickets

sold.

Box

$250 - 300 per ticket

Club Level

$150 - 190 per ticket

All other

From $70 to $110 per ticket

Attendance for the conference games was in the following ranges.

For all games except OU, the

number of tickets sold was the lower value. For the OU game, the maximum number of tickets were sold.

Box

1,400 – 2,000 tickets sold

Club Level

4,500 – 6,000 tickets sold

All other

45,000 – 50,000 tickets sold

Tolerable difference has been set at +/- 5% of the client’s recorded amount which is $27,235,240.

1.) What is your expectation for football ticket revenue for home games for the fiscal year-ending

6/30/2018?

2.) What is the difference between your expectation and the client’s recorded amount?

3.)What is the tolerable difference for this audit procedure? Do you need to do any further work?
Answer #1

1)

Non Conference Games ( 1 game) (150*1000) + (90*4000) + (50*40000) = 2510000

Conference Games ( 4 games) 4* [(250*1400) + (150*4500) + (70*45000) = 16700000

Conference games( 1 against OU) (300*2000) + (190*6000) + (110*50000) = 7240000

Total expected revenue = 26450000

2)

Difference - 27235240 - 26450000 = 785240

3)

Tolerable difference

26450000 + 5% = 27772500

Actual is 27235240 which is within the tolerable limit. No need to do any further work.